VIBRATIONAL ANALYSIS AND VALENCE FORCE FIELD FOR NITROTOLUENES, DIMETHYLANI-LINES AND SOME SUBSTITUTED METHYLBENZENES

\underline{B. VENKATRAM REDDY}, Department of Physics, Kakatiya University, Warangal-506 009, A.P., India Email: bvreddy67@yahoo.com; JAI KISHAN OJHA, Department of Physics, Government Degree College, Mancherial-504 208, A.P., India; G. RAMANA RAO, Department of Physics, Varada Reddy College of Engineering, Ananthasagar, Warangal-506 371, A.P.,India.

The Fourier transform infrared (FTIR) and Raman spectra of 2-amino-4-nitro-toluene; 2-amino-5-nitrotoluene; 2,4-dimethylaniline; 2,5-dimethylaniline; 2,6-dimethylaniline; 1,2,4-trimethylbenzene; 1,3,5-trimethylbenzene and pentamethyl-benzene have been recorded in the range 4000-400 cm^{−1} and 4000-30 cm^{−1}, respectively. A normal coordinate analysis was carried out for both in-plane and out-of-plane vibrations of these molecules using an 81-parameter modified valence force field. The force constants were refined using 251 frequencies of eight molecules in the Overlay leastsquare technique. The reliability of force constants was tested by making zero-order calculations for both in-plane and out-of plane vibrations for five related molecules. The potential energy distribution (PED) and eigen vectors calculated in the process were used to make unambiguous vibrational assignment of all the fundamentals.